A high-performance car initially travelling 97.2 km/h came to a stop in just 2.9 s. The mass of the car and its contents was 1850 kg. What was the magnitude of the average net force (braking force) on the car while it decelerated?


1.7 x 104N


4.9 x 104N


1.8x 104N


5.9 x 104N



1
Expert's answer
2021-11-11T12:04:15-0500

Solution.

"v_0=97.2 km\/h=27m\/s;"

"t=2.9s;"

"v=0 m\/s;"

"m=1850 kg;"

"a=\dfrac{v-v_0}{t};"

"a=\dfrac{0-27}{2.9}=-9.31 m/s^2;"

"F=ma;"

"F=9.31\sdot 1850=1.7\sdot 10^4N;"

Answer: "1.7 \sdot 10^4N."